LABORATORY SPECTROSCOPY OF THE C_{7}H AND C_{8}H RADICALS

\underline{MICHAEL C. McCARTHY}, MICHAEL J. TRAVERS, AND P. THADDEUS, Center for Astrophysics and Division of Applied Sciences, Harvard University, Cambridge, MA 02138; CARL A. GOTTLIEB, Division of Applied Sciences, Harvard University, Cambridge MA, 02138.

The linear carbon chain radicals C_{7}H and C_{8}H have been detected in their ^{2}Π electronic ground states in an acetylene/argon discharge. For each radical, both spin-orbit components have been detected, precise values for the rotational, fine structure, and lambda-doubling constants determined, and the entire fine structurerotational spectrum calculated. The lowest rotational transitions of C_{7}H and C_{8}H were also observed in a pulsed discharge supersonic molecular beam source by Fourier transform microwave spectroscopy. A summary of results, including an analysis of the hyperfine structure will be presented.